# Francisco Calvillo LPSM, Sorbonne Université, 4 Place Jussieu, 75005 Paris, France Luc Devroye

### v1

@@ 1 Introduction
Today’s online platforms have a record number of items available for their users. In many cases presenting them as a catalogue is no longer a good option and can make it difficult for the user to find what they are interested in. Recommendation systems have become an essential tool to cope with the overload of information available on the web.
@@
To model this more realistic scenario, our paper proposes a two-choice recommendation system inspired by the same idea. We envision a scenario where users are presented with two independent recommendations, drawn from the same temporal recommendation mechanism and allowing users to select the most appealing option. The resulting model has some intriguing mathematical properties and the main goal of this paper is to analyze the model in order to understand the long-term behavior of such recommendation systems.
@@ 1.1 A two-choice recommendation model
Each topic $i\leq 0$ is assigned a preference value $U_{i}$ within the range $[0,1]$, where we assume that $(U_{n})_{n\leq 0}$ is a sequence of random variables independent of the sequences $(Z_{n})_{n\geq 1}$ and $(W_{n})_{n\geq 1}$. These preference values are user-dependent.
@@
$\left\{\begin{array}[]{ll}C_{n}=C_{n-Z_{n}}\text{ and }V_{n}=V_{n-Z_{n}}&\text{ if }V_{n-Z_{n}}<V_{n-W_{n}}\\ C_{n}=C_{n-W_{n}}\text{ and }V_{n}=V_{n-W_{n}}&\text{ otherwise }~{}.\end{array}\right.$
Thus, by identifying the colour $C_{n}$ as the topic recommended at time $n\geq 1$, this definition means that at the time instance $n\geq 1$, the user receives two recommendations (the topic $C_{n-Z_{n}}$, and the topic $C_{n-W_{n}}$), and chooses the one with the lowest (best) preference value.
@@
These assumptions correspond to natural scenarios and allow us to study long-term consistency of the recommendation system. In the first case, topics have a monotone preference with the most recent (corresponding to $i=0$) being the most preferred one. In the second case, older topics are preferred, while in the third case topics have a random order of preference. We say that the system is consistent if, on the long run, it offers near-optimal recommendations to the user in terms of the their preference. In Section 1.3 we rigorously define various notions of consistency, corresponding to the scenarios above. These definitions lead us to the study of the sequence $(V_{n})_{n\geq 1}$ by studying the set of leaves $\mathcal{L}_{n}$. We will pay particular attention to the three following random variables:
@@
$\bullet$ $M_{n}=\min\left\{i:i\in\mathcal{L}_{n}\right\}$ (the leftmost leaf in $\mathcal{L}_{n}$)
$\bullet$ $L_{n}=|\mathcal{L}_{n}|$ (the number of leaves reached from $n$).
Each of these random variables helps us understand the long-term behavior of $V_{n}$ in each of the three initial configurations described above. One would ideally hope that $V_{n}\to\inf_{i\leq 0}U_{i}$. In configuration (i) this is equivalent to $R_{n}\to 0$, in configuration (ii) it is equivalent to $M_{n}\to-\infty$, while in case (iii) to $L_{n}\to\infty$. However, one cannot expect to have $R_{n}\to 0$ in a general case. To introduce a more reasonable consistency criterion for case (i), we require boundedness of the sequence $(R_{n})_{n\geq 1}~{}.$
Figure 1: An illustration of the definitions of $M_{n}$, $R_{n}$, and $\mathcal{L}_{n}$.
@@
and $\text{and }S_{n}=\{n\}\cup S_{n-\min(Z_{n},W_{n})}$
for any $n\geq 1$, so that $T_{n}^{Z}\subset T_{n}$ (resp. $T_{n}^{W}\subset T_{n}$) is the chain of vertices that can be reached from $n$ by following only the $Z$-edges (resp. $W$-edges), and $S_{n}\subset T_{n}$ is the chain of vertices that can be reached from $n$ by following only the shortest edges.
Finally, for any $n\geq 1$, let $r_{n}=\mathbb{P}\left(0\in T_{n}^{Z}\right)=\mathbb{P}\left(0\in T_{n}^{W}\right)$ denote the probability that vertex 00 belongs to the $Z$-chain starting at vertex $n$. We set $r_{0}=1$ so that $r_{n}$ satisfies the recursion
$r_{n}=\sum_{i=1}^{n}q_{i}r_{n-i}~{}.$ (1.1)
@@ 1.2 Related work
The random forest in [6] appears as a subgraph of a random graph model that was previously studied by Hammond and Sheffield in [10]. Given a sequence $(Z_{n})_{n\in\mathbb{Z}}$ of independent random variables, identically distributed in $\{1,2,3,\dots\}$, they consider a graph with vertex set $\mathbb{Z}$ such that vertices $n,m\in\mathbb{Z}$, with $m<n$, are connected by an edge if and only if $m=n-Z_{n}$. One can obtain the random forest in [6] by removing all edges between any two vertices with a non-positive index. They show that the random graph has almost surely a unique component when the sum $\sum_{n=1}^{\infty}r_{n}^{2}$ converges, and the graph almost surely has infinitely many connected components when the sum diverges.
Other variants of the (single-choice) random subtractive process were studied, in different contexts, by Blath, Jochen, González, Kurt, and Spano [4], Chierichetti, Kumar, and Tomkins [8], Baccelli and Sodre [3], Baccelli, Haji-Mirsadeghi, and Khezeli [2], Baccelli, Haji-Mirsadeghi, and Khaniha [1], and Igelbrink and Wakolbinger [12].
@@ 1.3 Results
We say that the system is consistent in configuration (ii) if $M_{n}\to-\infty$; while it is consistent in configuration (iii) if $L_{n}\to\infty$. We differentiate between strong and weak consistency based on whether convergence occurs almost surely or in probability. In the case of configuration (i), we say that the system is weakly consistent if $(R_{n})_{n\geq 1}$ is a tight sequence. It is strongly consistent if the sequence $(R_{n})_{n\geq 1}$ is almost surely bounded.
@@
Pareto tails. Consider a distribution with Pareto tails, that is, $q_{n}=\Theta(1/n^{1+\alpha})$, with $\alpha>0$. This implies that $p_{n}=\Theta(1/n^{\alpha})$. When $\alpha>1$, we see that $\mathbb{E}Z<\infty$. The more interesting situation is when $\alpha\in(0,1]$. For $\alpha\in(1/2,1]$, we have $\mathbb{E}Z=\infty$, yet $\mathbb{E}\min(Z,W)<\infty$. Finally, for $\alpha\in(0,1/2]$, we have $\mathbb{E}\min(Z,W)=\infty$. For light-tailed $Z$, we recall from Broutin, Devroye, Lugosi, Oliveira [6] that $r_{n}\to 1/\mathbb{E}Z$. For moderate and heavy tails, however, we have $r_{n}\to 0$. When $\alpha\in(0,1)$, then $Z$ is in the domain of attraction of the extremal stable law with parameter $\alpha$, which itself has a tail distribution function that decays at the rate $1/n^{\alpha}$ (see, e.g., Ibragimov and Linnik [11], Zolotarev [15] or Malevich [13]). One can then deduce that $r_{n}$ decays at the rate $1/n^{1-\alpha}$. In particular, $\sum_{n}r_{n}^{2}<\infty$ when $\alpha\in(0,1/2)$, which roughly corresponds to the case of heavy tails.
Next we describe the main results of the paper. For a summary, see the table at the end of this section.
@@
When $\mathbb{E}Z<\infty$, time jumps are so short that $T_{n}$ can’t reach distant leaves, thus forcing the set $\mathcal{L}_{n}$ to be bounded. This case does not require much work, since one can quickly determine consistency or inconsistency in each of the three configurations just by looking at the sequence $(M_{n})_{n\geq 1}$, which happens to be bounded almost surely.
@@
Let $M_{\infty}=\inf_{n\in\mathbb{N}}M_{n}$. If $\mathbb{E}Z<\infty$, then $M_{\infty}$ is finite almost surely. Thus, the sequences $(R_{n})_{n\in\mathbb{N}},(M_{n})_{n\in\mathbb{N}}$ and $(L_{n})_{n\in\mathbb{N}}$ are almost surely bounded.
@@
$\mathbb{P}\left(M_{\infty}\leq-x\right)\leq 2\sum_{n=1}^{\infty}\mathbb{P}\left(n-Z_{n}\leq-x\right)=2\sum_{n=1}^{\infty}p_{n+x}~{}.$
Since $\mathbb{E}Z<\infty$, we know that $\sum_{n=1}^{\infty}p_{n+x}$ is finite and goes to 00 as $x\rightarrow+\infty$. Thus, by continuity of measure,
$\mathbb{P}\left(M_{\infty}=-\infty\right)=\lim_{x\rightarrow+\infty}\mathbb{P}\left(M_{\infty}\leq-x\right)=0~{}.$
Since $|M_{n}|\geq L_{n}$ and $|M_{n}|\geq|R_{n}|$, this also implies that $(L_{n})_{n\in\mathbb{N}}$ and $(R_{n})_{n\in\mathbb{N}}$ are almost surely bounded.
@@
Let $Z$, $W$, $(Z_{n})_{n\geq 1}$ and $(W_{n})_{n\geq 1}$ be i.i.d. random variables. Assume furthermore that $\mathbb{E}\min(Z,W)=\infty~{}.$ Then, with probability one, the number of leaves satisfies
$\liminf_{n\to\infty}L_{n}<\infty~{}.$
In other words, the system is not strongly consistent in the heavy-tailed regime.
Proof.   Note that if $\min(Z_{n},W_{n})\geq n$, then $L_{n}\in\{1,2\}$. Thus, to prove Theorem 3, it suffices to show that with probability one, $\min(Z_{n},W_{n})\geq n$ happens infinitely often. This follows by the second Borel-Cantelli lemma since the events
($(\{\min(Z_{n},W_{n})\geq n\})_{n\in\mathbb{N}}$
@@
$\sum_{n\geq 1}\mathbb{P}\left(\min(Z_{n},W_{n})\geq n\right)=\mathbb{E}\left(\min(Z,W)\right)=\infty~{}.$
The behavior of the set $\mathcal{L}_{n}$ for moderate and heavy tails is better understood by examining its extreme points $M_{n}$ and $R_{n}$. The following two theorems are proved in Section 3.
@@
Given that the leftmost leaf in $\mathcal{L}_{n}$ goes to $\infty$- ∞ and the rightmost leaf remains bounded in the moderate heavy-tail regime (i.e., when $\mathbb{E}Z=\infty$ and $\mathbb{E}\min(Z,W)<\infty$), one can expect that the total number of leaves reached by the vertex $n$ diverges under these two assumptions. Thus, strong consistency is observed in each of the three configurations (i), (ii), and (iii).
On the other hand, only configuration (i) leads to strong consistency when the distribution of $Z$ has a heavy-tail (i.e., when $\mathbb{E}\min(Z,W)=\infty$). However, when $\sum_{n=1}^{\infty}r_{n}^{2}<\infty$, the model remains weakly consistent in configuration (iii), as established by the next two results. Theorems 6 and 7 are proved in Sections 4.1 and 4.2, respectively.
@@
Assume that the distribution of $Z$ is not supported by any proper additive subgroup of the integers and that $Z$ exhibits a moderate-sized tail (i.e., $\mathbb{E}\min(Z,W)<\infty$ and $\mathbb{E}Z=\infty$). Then
$L_{n}\to\infty\text{ almost surely. }$.
@@
When $Z$ has a heavy tail (i.e., $\mathbb{E}\min(Z,W)=\infty$), and $\sum_{n\geq 0}r_{n}^{2}<\infty~{},$ then
$L_{n}\to\infty\textit{ in probability.}$.
@@ 1.4 Summary
The following table summarizes our findings. Observe that weak and strong behavior coincide in most cases. They only differ in the extreme heavy tail case ($\mathbb{E}\min(Z,W)=\infty$). For an optimal user experience, one would need $L_{n}\to\infty$ almost surely, and the only case that assures this is when we have moderate tails. In addition, such moderate tails also guarantee strong consistency for the decreasing and increasing input scenarios (as is apparent from the strong consistency of $M_{n}$ and $R_{n}$).
@@
$M_{n}\to-\infty$ a.s. no (Thm. 2) yes (Thm. 4) yes (Thm. 4)
$L_{n}\to\infty$ a.s. no (Thm. 2) yes (Thm. 6) no (Thm. 3)
$R_{n}$ bounded a.s. yes (Thm. 2) yes (Thm. 5) no (Thm. 3)
@@
$M_{n}\overset{\mathbb{P}}{\to}-\infty$ no (Thm. 2) yes (Thm. 4) yes (Thm. 4)
$L_{n}\overset{\mathbb{P}}{\to}\infty$ no (Thm. 2) yes (Prop. 11) yes if $\sum r_{n}^{2}<\infty$ (Thm. 7)

### v2

# Subtractive random forests with two choices

@@ Abstract
Keywords: subtractive random forests, multi-choice recommendation
@@ 1 Introduction
Today’s online platforms have a record number of items available for their users. In many cases, presenting them as a catalogue is no longer a good option and can make it difficult for the user to find what they are interested in. Recommendation systems have become an essential tool to cope with the overload of information available on the web.
@@
To model this more realistic scenario, our paper proposes a two-choice recommendation system inspired by the same idea. We envision a scenario where users are presented with two independent recommendations, drawn from the same temporal recommendation mechanism and allowing users to select the most appealing option. The resulting model has some intriguing mathematical properties, and the main goal of this paper is to analyze the model in order to understand the long-term behavior of such recommendation systems.
While our model provides a mathematically tractable framework to study sequential recommendation dynamics, it is intentionally abstract and does not incorporate key aspects of modern recommender systems such as content information, user-specific features, or contextual data. This abstraction limits its direct applicability to real-world platforms, where personalization and diversity are central goals. However, we view this simplification as a necessary step: by first understanding the behavior of a minimal, analytically transparent model, we lay the groundwork for tackling more complex and realistic settings in which such additional factors can be systematically incorporated.
Beyond serving as a recommender system in its own right, our model can also be viewed as a theoretical tool to gain insight into more complex algorithms. In particular, it offers a simple framework to reason about phenomena such as the repeated occurrence of topics in classical recommender systems, where machine learning mechanisms are often treated as black boxes. By studying this simplified yet transparent model, we can uncover structural patterns that would otherwise remain hidden, offering insight into the dynamics of more sophisticated recommendation methods.
@@ 1.1 A two-choice recommendation model
Each topic $i\leq 0$ is assigned a preference value $U_{i}$ within the range $[0,1]$, where we assume that $(U_{n})_{n\leq 0}$ is a sequence of random variables independent of the sequences $(Z_{n})_{n\geq 1}$ and $(W_{n})_{n\geq 1}$. In the context of our model, the term user refers to a representative individual interacting with the system, and the values $U_{i}$ represent how appealing each topic is to this user.
@@
[…]
Thus, by identifying the colour $C_{n}$ as the topic recommended at time $n\geq 1$, this definition means that at the time instant $n\geq 1$, the user receives two recommendations (the topic $C_{n-Z_{n}}$, and the topic $C_{n-W_{n}}$), and chooses the one with the lowest (best) preference value.
@@
These assumptions correspond to natural scenarios and allow us to study the long-term consistency of the recommendation system. In the first case, topics have a monotone preference with the most recent (corresponding to $i=0$) being the most preferred one. In the second case, older topics are preferred, while in the third case, topics have a random order of preference. We say that the system is consistent if, in the long run, it offers near-optimal recommendations to the user in terms of their preference. In Section 1.3, we rigorously define various notions of consistency, corresponding to the scenarios above. These definitions lead us to study the sequence $(V_{n})_{n\geq 1}$ by studying the set of leaves $\mathcal{L}_{n}$. We will pay particular attention to the three following random variables:
@@
$\bullet$ $L_{n}=\min\left\{i:i\in\mathcal{L}_{n}\right\}$ (the leftmost leaf in $\mathcal{L}_{n}$)
$\bullet$ $_{n}=|\mathcal{L}_{n}|$ (the number of leaves reached from $n$).
Each random variable helps us understand the long-term behavior of $V_{n}$ in the three initial configurations described above. Ideally, one would hope that $V_{n}\to\inf_{i\leq 0}U_{i}$. In configuration (i) this is equivalent to $R_{n}\to 0$, in configuration (ii) it is equivalent to $L_{n}\to$-$\infty$, while in case (iii) to $_{n}\to\infty$. However, one cannot expect to have $R_{n}\to 0$ in a general case. To introduce a more reasonable consistency criterion for case (i), we require boundedness of the sequence $(R_{n})_{n\geq 1}~.$
@@
$\text{and }S_{n}=\{n\}\cup S_{n-\min(Z_{n},W_{n})}$
for any $n\geq 1$, so that $T_{n}^{Z}\subset T_{n}$ (resp. $T_{n}^{W}\subset T_{n}$) is the set of vertices that can be reached from $n$ by following only the $Z$-edges (resp. $W$-edges), and $S_{n}\subset T_{n}$ is the set of vertices that can be reached from $n$ by following only the shortest edges.
Figure 1: Illustration of the subgraph induced by $T_{n}$. Shown are the definitions of $L_{n}$, $R_{n}$, and $\mathcal{L}_{n}$. Blue edges indicate the $Z$-chain of vertices $T_{n}^{Z}$, while red edges connect all vertices in the chain $S_{n}$.
Finally, for any $n\geq 1$, let $r_{n}=\mathbb{P}\left(0\in T_{n}^{Z}\right)=\mathbb{P}\left(0\in T_{n}^{W}\right)$ denote the probability that vertex 0 belongs to the $Z$-chain starting at vertex $n$. We set $r_{0}=1$ so that $r_{n}$ satisfies the recursion
$r_{n}=\sumop\displaylimits_{i=1}^{n}q_{i}r_{n-i}~.$ (1.1)
For convenience, a table summarizing the main notation used throughout the paper is provided in the Appendix.
Remark 1.
Observe that the chains $T_{n}^{Z}$, $T_{n}^{W}$, and $S_{n}$ can be naturally interpreted as a renewal process, where each step along the chain corresponds to a renewal event, and the edge lengths play the role of renewal intervals. Within this framework, the quantity $r_{n}$ coincides with the renewal measure of the process corresponding to $T_{n}^{Z}$.
@@ 1.2 Related work
The random forest in [6] appears as a subgraph of a random graph model previously studied by Hammond and Sheffield in [10]. Given a sequence $(Z_{n})_{n\in\mathbb{Z}}$ of independent random variables, identically distributed in $\{1,2,3,\dots\}$, they consider a graph with vertex set $\mathbb{Z}$ such that vertices $n,m\in\mathbb{Z}$, with $m<n$, are connected by an edge if and only if $m=n-Z_{n}$. One can obtain the random forest in [6] by removing all edges between any two vertices with a non-positive index. They show that the random graph has almost surely a unique component when the sum $\sumop\displaylimits_{n=1}^{\infty}r_{n}^{2}$ converges, and the graph almost surely has infinitely many connected components when the sum diverges.
Other variants of the (single-choice) random subtractive process have been analyzed in related probabilistic models. Blath, González, Kurt, and Spano [4] studied long-range seed bank models, where the genealogical structure of populations is governed by random time jumps analogous to the subtractive mechanism. Chierichetti, Kumar, and Tomkins [8] examined sequence models whose asymptotic behavior shares similarities with renewal-type processes underlying our setting. Baccelli, Haji-Mirsadeghi, Khezeli, and Sodre [3, 2, 1] explored subtractive dynamics in the context of unimodular random graphs and population processes. Finally, Igelbrink and Wakolbinger [12] investigated reinforced urn schemes where coalescence probabilities play a role closely related to recurrence properties of subtractive chains.
@@ 1.3 Results
We say that the system is consistent in configuration (ii) if $L_{n}\to$-$\infty$; while it is consistent in configuration (iii) if $_{n}\to\infty$. We differentiate between strong and weak consistency based on whether convergence occurs almost surely or in probability. In the case of configuration (i), we say that the system is weakly consistent if $(R_{n})_{n\geq 1}$ is a tight sequence. It is strongly consistent if the sequence $(R_{n})_{n\geq 1}$ is almost surely bounded.
@@
[…]
Next, we describe the main results of the paper. For a summary, see the table at the end of this section.
@@
When $\mathbb{E}Z<\infty$, time jumps are so short that $T_{n}$ can’t reach distant leaves, thus forcing the set $\mathcal{L}_{n}$ to be bounded. This case does not require much work, since one can quickly determine consistency or inconsistency in each of the three configurations just by looking at the sequence $(L_{n})_{n\geq 1}$, which happens to be bounded almost surely.
@@
Let $L_{\infty}=\inf_{n\in\mathbb{N}}L_{n}$. If $\mathbb{E}Z<\infty$, then $L_{\infty}$ is finite almost surely. Thus, the sequences $(R_{n})_{n\in\mathbb{N}},(L_{n})_{n\in\mathbb{N}}$ and $({}_{n})_{n\in\mathbb{N}}$ are almost surely bounded.
@@
$\mathbb{P}\left(L_{\infty}\leq-x\right)\leq 2\sumop\displaylimits_{n=1}^{\infty}\mathbb{P}\left(n-Z_{n}\leq-x\right)=2\sumop\displaylimits_{n=1}^{\infty}p_{n+x}~.$
Since $\mathbb{E}Z<\infty$, we know that $\sumop\displaylimits_{n=1}^{\infty}p_{n+x}$ is finite and goes to 0 as $x\rightarrow+\infty$. Thus, by continuity of measure,
$\mathbb{P}\left(L_{\infty}=-\infty\right)=\lim_{x\rightarrow+\infty}\mathbb{P}\left(L_{\infty}\leq-x\right)=0~.$
Since $|L_{n}|\geq{}_{n}$ and $|L_{n}|\geq|R_{n}|$, this also implies that $({}_{n})_{n\in\mathbb{N}}$ and $(R_{n})_{n\in\mathbb{N}}$ are almost surely bounded.
@@
Let $Z$, $W$, $(Z_{n})_{n\geq 1}$ and $(W_{n})_{n\geq 1}$ be i.i.d. random variables. Assume furthermore that $\mathbb{E}\min(Z,W)=\infty~.$ Then, with probability one,
$\liminf_{n\to\infty}{}_{n}$$\infty~$.
In other words, the system is not strongly consistent in the heavy-tailed regime in configuration (iii).
Proof. Note that if $\min(Z_{n},W_{n})\geq n$, then $_{n}\in\{1,2\}$. Thus, to prove Theorem 3, it suffices to show that with probability one, $\min(Z_{n},W_{n})\geq n$ happens infinitely often. This follows by the second Borel-Cantelli lemma since the events
$(\{\min(Z_{n},W_{n})\geq n\})_{n\geq 0}$
@@
$\sumop\displaylimits_{n\geq 0}\mathbb{P}\left(\min(Z_{n},W_{n})\geq n\right)=\mathbb{E}\left(\min(Z,W)\right)=\infty~.$
The behavior of the set $\mathcal{L}_{n}$ for moderate and heavy tails is better understood by examining its extreme points $L_{n}$ and $R_{n}$. The following two theorems are proved in Section 3.
@@
Given that the leftmost leaf in $\mathcal{L}_{n}$ goes to -$\infty$ and the rightmost leaf remains bounded in the moderate heavy-tail regime (i.e., when $\mathbb{E}Z=\infty$ and $\mathbb{E}\min(Z,W)<\infty$), one can expect that the total number of leaves reached by the vertex $n$ diverges under these two assumptions. Thus, strong consistency is observed in each of the three configurations (i), (ii), and (iii).
On the other hand, only configuration (i) leads to strong consistency when the distribution of $Z$ has a heavy-tail (i.e., when $\mathbb{E}\min(Z,W)=\infty$). However, when $\sumop\displaylimits_{n=1}^{\infty}r_{n}^{2}<\infty$, the model remains weakly consistent in configuration (iii), as established by the next two results. Theorems 6 and 7 are proved in Sections 4.1 and 4.2, respectively.
@@
Assume that the distribution of $Z$ is aperiodic, that is, the greatest common divisor of the support of $Z$ is $1$. If $Z$ exhibits a moderate-sized tail (i.e., $\mathbb{E}\min(Z,W)<\infty$ and $\mathbb{E}Z=\infty$), then
$_{n}\to\infty\text{ almost surely. }$
@@
When $Z$ has a heavy tail (i.e., $\mathbb{E}\min(Z,W)=\infty$), and $\sumop\displaylimits_{n\geq 0}r_{n}^{2}<\infty~,$ then
$_{n}\to\infty\textit{ in probability.}$
@@ 1.4 Summary
The following table summarizes our findings. Observe that weak and strong behavior coincide in most cases. They only differ in the extreme heavy tail case ($\mathbb{E}\min(Z,W)=\infty$). For an optimal user experience, one would need $_{n}\to\infty$ almost surely, and the only case that assures this is when we have moderate tails. In addition, such moderate tails also guarantee strong consistency for the decreasing and increasing input scenarios (as is apparent from the strong consistency of $L_{n}$ and $R_{n}$).
@@
$L_{n}\to$-$\infty$ a.s. no (Thm. 2) yes (Thm. 4) yes (Thm. 4)
$_{n}\to\infty$ a.s. no (Thm. 2) yes (Thm. 6) no (Thm. 3)
$R_{n}$ bounded a.s. yes (Thm. 2) yes (Thm. 5) no (Thm. 5)
@@
$L_{n}\overset{\mathbb{P}}{\to}$-$\infty$ no (Thm. 2) yes (Thm. 4) yes (Thm. 4)
$_{n}\overset{\mathbb{P}}{\to}\infty$ no (Thm. 2) yes (Prop. 11) yes if $\sumop\displaylimits r_{n}^{2}<\infty$ (Thm. 7)
@@
1.5 Simulations
To complement our theoretical analysis, we conducted a series of simulations to examine the empirical behavior of the model. We considered discrete Pareto distributions for the random variables $Z$ and $W$, with probabilities $q_{n}=c/n^{1+\alpha}$, so that varying $\alpha$ produces light-tailed, moderate-tailed, and heavy-tailed regimes. Our analysis focused on the evolution of the preferential value $V_{n}$ of the topic recommended at time $n$ in configuration (iii), where topics are assigned independent uniform preference values.
Figure 3 presents single realizations of $V_{n}$ for each tail regime, illustrating the emergence (or absence) of strong consistency. Figure 3 shows the corresponding averages over $1000$ independent simulations, which emphasize the form of weak consistency predicted in the heavy-tailed setting. The simulations reveal three qualitatively distinct behaviors:
  • Light-tailed regime ($\alpha=1.2$): no evidence of consistency; recommendations stabilize around a single topic, and $V_{n}$ does not converge to 0.
  • Moderate-tailed regime ($\alpha=0.6$): strong consistency, with the process converging reliably to a dominant topic.
  • Heavy-tailed regime ($\alpha=0.3$): strong inconsistency at the trajectory level, but weak consistency visible in the averaged dynamics, in agreement with Remark 2.
Figure 2: Simulation of the evolution of $V_{n}$ in configuration (iii) for a single run across light-tailed, moderate-tailed, and heavy-tailed regimes.
Figure 3: Evolution of the empirical mean of $V_{n}$ in configuration (iii) over $1000$ independent simulations for light-tailed, moderate-tailed, and heavy-tailed regimes.
Configurations $(i)$ and $(ii)$ were also examined, and the corresponding simulation outcomes are in full agreement with the theoretical results established for these regimes. Since their qualitative behavior is straightforward and consistent with the analytical predictions, the corresponding plots are omitted for brevity.
In addition to tracking the evolution of $V_{n}$, it is also instructive to visualize the subgraph induced by the vertices $T_{n}$, consisting of the vertices reached by vertex $n$. Coloring vertices with positive index in blue and vertices with non-positive index (leaves or topics) in red provides a structural perspective on the process. The resulting plots (see Figure 4) make the differences between the three regimes strikingly clear: in the light-tailed case, the short edges yield a dense graph with many blue vertices but very few leaves; in the moderate-tailed case, the number of leaves grows substantially while blue vertices remain numerous; and in the heavy-tailed case, the dominance of long edges produces a sparse structure with very few blue vertices and a predominance of leaves. These observations are fully consistent with our theoretical analysis and complement the intuition provided by the $V_{n}$ plots.
Figure 4: Visualization of the subgraph induced by $T_{n}$ in the light-, moderate-, and heavy-tailed regimes, with $n=150$.